Albert is 19 years older than Pat. The sum of their ages is 31. How old are they?
Found 2 solutions by TimothyLamb, algebrahouse.com:
Answer by TimothyLamb(4379)

a = p + 19
a + p = 31
---
a + p = 31
p + 19 + p = 31
2p = 12
---
p = 6
a = 25
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by algebrahouse.com(1659)

x = Pat's age
x + 19 = Albert's age {Albert is 19 years older than Pat}

x + x + 19 = 31 {sum of their ages is 31}
2x + 19 = 31 {combined like terms}
2x = 12 {subtracted 19 from each side}
x = 6 {divided each side by 2}
x + 19 = 25 {substituted 6, in for x, into x + 19}

Pat is 6 and Albert is 25

For more help from me, visit: www.algebrahouse.com